240 people running a race 160 of them  have on white shirts 130 of them have on white shorts 68 of them have on neither how many have on both?
Answer by alicealc(293)

Assume that number of people have both white shirts and white shorts are x

160 - x + 130 - x + x + 68 = 240
358 - x = 240
358 - 240 = x
118 = x

so, there are 118 people have both white shirts and white shorts